

********************************************************************
UARS WEEKLY STATUS REPORT - 05 MAY 2000
********************************************************************
***********
SOLSTICE
***********
We continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations.
SOLSTICE operates normally, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. Level 2 processing is halted due to bugs in the OA_SSPP_ATT
service. CDHF and CSC personnel are working on the problem.
***********
SUSIM
***********
SUSIM continues to gather a near normal set of daily solar scans
scheduled for periods of predicted telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
HALOE continues successful operations with all engineering parameters
within nominal operating region. HALOE will be powered off some time May
7th as the spacecraft enters the graze period.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG participated in the Monthly Status Review Meeting, held on 3
May. The MPG distributed the Instrument Operations Schedule for the next
yaw period: 4 June to 4 July 2000.
All the systems in the MPG room were shutdown on May 4 for the power
reconfiguration operation. All the systems are operational after reboot.
Routine processing of instrument Daily Activity Plans continues with no
problems.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
- All requested reprocessing is complete as of this reporting period
The CDHF will be sending out an updated version of the NASA Security
Rules of Behavior to all users, in accordance with NASA ISO requirements.
UARS CDHF REHOST:
- Successful processing on Rehost System:
Attitude Extraction, Observatory Log, Sensor Extraction, HALOE, MLS,
SUSIM, Solstice, and WINDII.
- An electronic ingest of LEVEL0, UARS DAY ,3144,3155,3156,3157,3158 was
received from the LZP. Validation in progress.
Sensor Extraction Software was processed for 3144. FDF has performed
a successful quality check against the Sensor Files.
- Test HRDI V4.4 on the Rehost System.
- Preparing to test ISAMS V12.0 and PEM V3.32 software on the
Rehost System.
- Preparing to test DAAC transfer software.
***********
FOT
***********
REPORTING PERIOD
This report covers 22 April, 2000 (Orbit 47073, GMT Day 113) through 28 April
2000 (Orbit 47177, GMT Day 119).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 113/0408 47075 Off
HALOE 119/1152 47170 On
HALOE 119/2130 47176 Off
HALOE 119/2130 47176 On
A Reverse-to-Forward Yaw Around maneuver (number 88) was successfully executed
in Orbit 47127 beginning 116/15:52:58 and completing at 116/15:30:08 on 25
April 2000. The SSPP Instruments (ACRIM, SOLSTICE, SUSIM) were enabled as well
as HRDI and PEM in day only operations. Due to the SA position at 269 degrees,
the effective night length of the spacecraft after the maneuver was
approximately 51 minutes. Post maneuver analysis of data indicate that during
the Yaw Around, the end-of-night load bus voltage was 23.5 volts and the
average depths of discharge (DODs) on Batteries 2 and 3 were approximately 22.0
and 22.5 percent respectively. The next (forward-to-reverse) YAW is scheduled
for 05 June 2000 (DOY 157).
The beta angle ranged between the angles of 17.7 degrees to 12.5 degrees for
this report period. The beta angle is now increasing toward a maximum angle of
+76.0 degrees on 15 May 2000 (DOY 136).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.3 degrees C. Additional battery performance data for each
battery on 21 April 2000 and 28 April 2000 are provided in Appendix D.
The clock error ranged from 3.4 to -9.8 msec during this report period. Clock
rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 07
% ALERTS = 01
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 06
% HALOE = 05
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 03
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
- Transponder A: 00/091 - 2287.496948
- Transponder B: 00/091 - 2287.499113
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/118 show the observatory data loss to be 295 hours, 29
minutes, 41.209 seconds (an increase of 23 minutes 12.064 seconds since the
last report period).
The increase of 23 minutes 12.064 seconds occurred over one (1) day:
- 23 minutes 12.064 seconds on Day 116 due to POCC scheduling error.
These calculations have been recomputed to include 1999/346 through 2000/006
statistics using scheduled times.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to have problems properly loading the new TMONs to the
simulator. Both the FOT and its support staff are doing everything possible to
identify and correct this problem so that the testing of the new safehold TMONs
can begin.
The FOT continues to develop an online procedure for safehold should it happen
before the new TMONs are in place or during and OBC halt. This procedure has
yet to be finalized and must undergo peer review before it will be ready.
MEETINGS
In support of the UARS Upgrade task, the FOT attended a meeting to discuss the
basic roles and activities of each FOT member.
The FOT held a meeting with NASA battery experts on 28 April 2000. The goal of
the meeting was to determine the proper offset angle for the solar array in the
new online procedure for safehold. Identifying this angle will require more
analysis and will hopefully be determined sometime next week.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for June 05, 2000 (DOY 157). This maneuver
will be a Forward to Reverse Flight maneuver (Northern Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 28 April 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
47074 113/0231 00-173 NBTR NBTR PLAYBACK
47075 113/0408 00-174 HALOE EARLY OFF- S/C LOW PWR
47075 113/0453 00-173 NBTR NBTR OPS DAY 113
47089 114/0314 00-175 NBTR NBTR PLAYBACK
47167 119/0749 00-176 NBTR NBTR OPS DAY 119
47170 119/1152 00-177 HALOE Power ON
47170 119/1249 00-178 OBC Clk Adj 0.00 to -20.6 msec/day
47174 119/1806 00-179 NBTR NBTR OPS DAY 120
47174 119/1837 00-180 HALOE Power ON
47176 119/2130 00-181 HALOE Power OFF
47176 119/2130 00-182 HALOE Power ON
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 28 April 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPSEQBIMON Y-HI PSE Quiet Bus Current Mon 47074
HRPSEQBIMON Y-HI PSE Quiet Bus Current Mon 47105-8
HRPTSPBIMON R-LO PTS Pulse Bus Current Mon 47117
HRPTSPBIMON R-LO PTS Pulse Bus Current Mon 47117-8
HRPSEQBIMON Y-HI PSE Quiet Bus Current Mon 47120-22
OBSVLBEON Y-LO Load Bus Volts at EON 47129
HRPTSPBIMON R-LO PTS Pulse Bus Current Mon 47135-8
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
OBSZ1OOL UFLTR Z1 Residual Exceeded 100 Arcsec 47127
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 28 April 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
47170 119/12:51 (28 Apr 00) from 0.00 to -20.6 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 28 April 2000
21 April 2000 - GMT Day 112 Beta = 17.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ----------- ------ ------- ------
1 -22.4/-44.8 -3.05/-4.31 +0.0/+0.0 15.8 0.000 0.0
2 +84.0/-5.6 +6.44/+5.15 +30.4/-14.4 24.0 1.030 22.5
3 +33.6/-44.8 +5.15/+3.87 +29.6/-15.2 23.8 1.030 22.0
28 April 2000 - GMT Day 119 Beta = 12.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ---------- ------- ------- ------
1 -22.4/-44.8 -3.68/-4.94 +0.0/+0.0 15.8 0.000 0.0
2 +33.6/-11.2 +5.47/+3.87 +30.4/-14.4 24.3 1.022 20.5
3 +33.6/-5.6 +4.51/+2.60 +29.6/-15.2 24.2 1.020 20.1
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ----- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 28 April 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-016 FOT 47130 116/2012 22279 HGA slew error
00-018 FOT 47165 119/0415Z N/A AP-8 Crashed
00-019 FOT 47166 119/0635 N/A AP-8 Crashed
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-016 FOT 47130 116/2012 2227 HGA slew error
00-018 FOT 47165 119/0415Z N/A AP-8 Crashed
00-019 FOT 47166 119/0635 N/A AP-8 Crashed
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-002 FOT 47130 116/2012 22279 HGA slew error
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiv 512
00-007 FOT see list April 00 14170 Generic Late Acq.
SEVNTFW00/17